OpenStudy (anonymous):

Working his way through school, Joe works two part-time jobs for a total of 36 hours a week. Job A pays $5.70 per hour, and Job B pays $7.10 per hour. How many hours did he work at each job the week that he made $230.40. (round to two decimal places if necessary).

OpenStudy (anonymous):

\[5.7x+7.1(36-x)=230.4\] guy needs a new job

OpenStudy (anonymous):

solve to get 18 for x, so he worked 18 hours in each place

OpenStudy (anonymous):

i do need a new job =/

OpenStudy (anonymous):

hope the solution is clear. easiest to clear the decimals and write \[57x+71(36-x)=2304\] \[57x+255-71x=2304\] \[-14x=-252\] \[x=\frac{252}{14}=18\]

OpenStudy (anonymous):

so its 18 hours for Job A and Job B right?

OpenStudy (anonymous):

yeah thats what it looks like :) since the total is 36 hours.